# AFIRE: Accurate and Fast Image Reconstruction Algorithm for Geometric-inconsistency Multispectral CT

### v1

@@ Abstract
For nonlinear multispectral computed tomography (CT), accurate and fast image reconstruction is challenging when the scanning geometries under different X-ray energy spectra are inconsistent or mismatched. Motivated by this, we propose an accurate and fast algorithm named AFIRE to address such problem in the case of mildly full scan. We discover that the derivative operator (gradient) of the involved nonlinear mapping at some special points, for example, at zero, can be represented as a composition (block multiplication) of a diagonal operator (matrix) composed of X-ray transforms (projection matrices) and a very small-scale matrix. Based on the insights, the AFIRE is proposed respectively from the continuous, discrete and actual-use perspectives by leveraging the simplified Newton method. Under proper conditions, we establish the convergence theory of the proposed algorithm. Furthermore, numerical experiments are also carried out to verify that the proposed algorithm can accurately and effectively reconstruct the basis images in completely geometric-inconsistency dual-energy CT with noiseless and noisy projection data. Particularly, the proposed algorithm significantly outperforms some state-of-the-art methods in terms of accuracy and efficiency. Finally, the flexibility and extensibility of the proposed algorithm are also demonstrated.
@@
Multispectral CT, dual-energy CT, geometric inconsistency, large-scale nonlinear inverse problem, accurate and fast image reconstruction algorithm, one-step method, convergence analysis
@@ 1 Introduction
Basically, multispectral computed tomography (MSCT) is an imaging modality aiming to reconstruct material-specific (basis) images from multiple sets of attenuation data collected with different X-ray energy spectra, which can subsequently be used to form virtual monochromatic images (VMIs) or the spatial distributions of linear attenuation coefficient (LAC) at energies of interest by linear combination [18, 3, 27]. MSCT has attracted increasing attention and research interest due to the development of hardware and the requirement of applications, as well as its advantages in considering the X-ray energy-dependent information over the conventional CT such as artifact suppression, contrast improvement and object separation [27]. Dual-energy CT (DECT), as the first generation of MSCT, has been implemented in clinical practice for about twenty years [17]. Despite the long history of DECT, its clinical and technical research is still in its infancy [19].
@@
It is well-known that the forward data model used in conventional CT assumes that all of the X-ray photons emitting from an X-ray source are monochromatic, namely, have the same energy [32]. This assumption makes the model linear, which causes the image reconstruction to become a linear inverse problem, but leading to difficulties in artifact suppression and limitations in material differentiation. In real-world CT, however, this assumption is not realistic, since the output X-ray beams are polychromatic, which consist of photons with a wide range of energies that distributed as the energy spectrum. The LAC of the scanned material depends not only on its own elemental composition but also on the energy of the probing X-ray photons. Hence, the forward data model of MSCT may be described as a nonlinear mapping, and the corresponding image reconstruction can be formulated as a nonlinear inverse problem [3, 26, 29, 19, 27, 21].
@@
Although development of algorithms for reconstructing basis images in MSCT constitutes an important topic of active research, accurate and fast image reconstruction remains a challenging task, largely because its forward data model is nonlinear. In this work, we understand different reconstruction algorithms from the perspective of whether the scanning geometric parameters are consistent or matched under different X-ray energy spectra.
@@
One situation is that the scanning geometric parameters are consistent under different energy spectra. That is to say, for each X-ray path, under every energy spectrum, there is always an X-ray beam passing through it. Here this case is called geometric-consistency MSCT, in which the image reconstruction can be accurately realized by a two-step method: step (i) estimates the basis sinograms from measurements through solving a lot of very small-scale independent systems of nonlinear equations, whereas step (ii) reconstructs basis images from the basis sinograms estimated. This method is commonly referred to as two-step data-domain-decomposition (DDD) method. As step (ii) involves the inversion of a straightforward linear system as in conventional CT, step (i) constitutes the key step of the DDD method in which a nonlinear data model needs to be inverted. Although many algorithms can invert the nonlinear model accurately, analysis of its solution existence, uniqueness and stability remains difficult. In recent years, several works studied the solution properties such as [23, 2, 4, 11, 16].
@@
The other is that the scanning geometric parameters are inconsistent or mismatched under different energy spectra. Namely, there is at least an X-ray path that all the X-ray beams under some energy spectrum do not pass through it. We call this case the geometric-inconsistency MSCT. A more extreme example is that for every X-ray path, there is always an energy spectrum under which no X-ray beams pass through it, which can be seen as completely geometric-inconsistency MSCT. The configurations of many commercial devices fall into this situation, such as the fast kV-switching DECT of GE [27].
@@
The development of accurate and fast reconstruction algorithms for this case is more challenging than the geometric-consistency case. The authors in [35] matched the originally mismatched neighboring projections of high- and low-energy spectra in DECT, and then proposed a two-step DDD method to invert those one-view-deviated projection pairs. Similarly, using the matched projections as in [35], an iterative filtered back-projection (FBP) reconstruction method was proposed in [24], which applies Newton’s method with only once iteration and FBP alternately. However, due to the different geometric parameters of the matched projections, these approaches cannot result in accurate image reconstruction. In the last two decades, progress has been made in developing one-step methods for accurate image reconstruction in MSCT. To reconstruct the basis images directly, such kind of method often needs to devise an optimization model from a certain perspective (e.g., statistics) and then develops an iterative scheme to solve that model [13, 6, 33, 25, 5, 7, 28, 8, 15], or considers the problem as solving a large-scale system of nonlinear equations [34, 14]. Because of the nonlinear data model in MSCT, the optimization models designed are generally non-convex. While these methods can often achieve high image quality through a large number of iterations, they are time-consuming, and their convergence analyses were mostly not considered.
@@
In this article, we aim to develop an image reconstruction algorithm of the one-step type that takes into account both accuracy and efficiency, has mathematical guarantee, and is suitable for the geometric-inconsistency MSCT.
@@
The rest of this paper is organized as follows. In section 2, we introduce the forward data models of MSCT. In section 3, we propose a novel algorithm to solve the image reconstruction problem in geometric-inconsistency MSCT. We then prove the convergence of the proposed algorithm in section 4. In section 5, numerical experiments are carried out to illustrate the performance of the proposed algorithm. The flexibility and extensibility of the proposed algorithm are discussed in section 6. Finally, we conclude the paper in section 7.
@@ 2.1 CC-data model
and $\mu(\boldsymbol{y},E)$ represents the energy-dependent LAC of interest at spatial position $\boldsymbol{y}\in\mathbb{R}^{n}$ ($n=2$ or $3$) and energy $E\in[0,E_{\text{max}}]$, and $\mathbb{R}^{n}$ the real Euclidean space, $E_{\text{max}}$ the maximum energy in the scan. The function $\mu$ of LAC is decomposed often into a linear combination of basis images $\{f_{d}\}_{d=1}^{D}$ as
@@
where $b_{d}(E)$ denotes the expansion coefficient, and $D$ the number of basis images. For example, for the typical X-ray energy range in diagnostic DECT, $D=2$ is chosen often as photoelectric effect and Compton scatter are the dominant factors contributing to $\mu$; $b_{d}$ denotes the known function of mass attenuation coefficient (MAC) of the $d$-th basis material (e.g., water and bone); and $f_{d}$ represents the material-equivalent density function of the $d$-th basis material.
@@
The formula in 3 is referred to as the CC-data model because $\theta^{[q]}$, $E$, and $\boldsymbol{y}$ are continuous variables.
@@ 2.2 DD-data model
In practice, data can be detected only for discrete energies and rays, and an image is reconstructed only on an array of discrete pixels/voxels. Hence, it is possible to develop a DD-data model, which may be constructed by implementing a specific discrete scheme into the CC-data model in 3. However, in general, there is no direct connection between the two models.
Assume that the energy domain $[0,E_{\text{max}}]$ is divided into $M$ intervals with equal size $\Delta_{E}$; the data $g_{j}^{[q]}$ are detected for discrete directions $\theta_{j}^{[q]}$ with $1\leq j\leq J_{q}$ under the $q$-th spectrum; each basis image $f_{d}$ is discretized by an array of size $I$. Then a DD-data model is obtained as
$g_{j}^{[q]}=\ln\sum_{m=1}^{M}s^{[q]}_{m}\exp\left(-\sum_{d=1}^{D}b_{dm}\sum_{i=1}^{I}p^{[q]}_{ji}f_{di}\right)\quad\text{for $1\leq j\leq J_{q}$,}$ (4)
@@ 3.1.1 Continuous MSCT image reconstruction problem
For simplicity, define $\mathbb{T}^{n}:=\{(\boldsymbol{\omega},\boldsymbol{x})|\boldsymbol{\omega}\in\mathbb{S}^{n-1},\boldsymbol{x}\in\boldsymbol{\omega}^{\bot}\}$, and
$\mathcal{P}f(\boldsymbol{\omega},\boldsymbol{x}):=\int_{-\infty}^{+\infty}f(\boldsymbol{x}+t\boldsymbol{\omega})\mathrm{d}t,\quad(\boldsymbol{\omega},\boldsymbol{x})\in\mathbb{T}^{n},$ (5)
which is the X-ray transform of a certain function $f$ defined on $\mathbb{R}^{n}$, equivalently the Radon transform if $n=2$ [29]. Here $\mathbb{S}^{n-1}$ represents the unit sphere in $\mathbb{R}^{n}$, and $\boldsymbol{\omega}^{\bot}$ denotes the hyperplane perpendicular to $\boldsymbol{\omega}$.
As the X-ray $l(\theta^{[q]})$ in 3 can be parametrized by $(\boldsymbol{\omega},\boldsymbol{x})$, we rewrite 5 as
$\mathcal{P}f(\boldsymbol{\omega},\boldsymbol{x})=\int_{l(\theta^{[q]})}f(\boldsymbol{y})\mathrm{d}l.$ (6)
@@
$K^{[q]}(\boldsymbol{f}):=\ln\int_{0}^{E_{\max}}s^{[q]}(E)\exp\left(-\sum_{d=1}^{D}b_{d}(E)\mathcal{P}f_{d}\right)\mathrm{d}E,$ (7)
@@
Consequently, the aim of the continuous MSCT image reconstruction is to reconstruct $\boldsymbol{f}$ from the measured data functions $\{g^{[q]}\}_{q=1}^{Q}$ such that
@@
where $\boldsymbol{K}(\boldsymbol{f})=[K^{[1]}(\boldsymbol{f}),\ldots,K^{[Q]}(\boldsymbol{f})]^{\boldsymbol{\mathsf{T}}}$ and $\boldsymbol{g}=[g^{[1]},\ldots,g^{[Q]}]^{\boldsymbol{\mathsf{T}}}$.
@@ 3.1.2 The proposed algorithm for the continuous problem
Here we propose the algorithm AFIRE for solving the continuous MSCT image reconstruction problem.
Before that, we introduce the required notation. Let $\mathcal{X}$ be a normed vector space with norm $\|\cdot\|_{\mathcal{X}}$, and
$\mathcal{X}^{D}:=\underbrace{\mathcal{X}\times\cdots\times\mathcal{X}}_{D}$
be the Cartesian product of the space $\mathcal{X}$ with the norm defined by
$\|\boldsymbol{u}\|_{{\mathcal{X}}^{D}}:=\left(\sum_{d=1}^{D}\|u_{d}\|^{2}_{\mathcal{X}}\right)^{1/2},$
@@
Denote the X-ray transform in 5 by the mapping
$\mathcal{P}:\mathcal{F}=H_{0}^{-1/2}(\Omega)\longrightarrow\tilde{\mathcal{G}}=\mathcal{P}(\mathcal{F})\subset\mathcal{L}^{2}(\mathbb{T}^{n})=\mathcal{G}.$
By [30, theorem 2.18], it is known that $\mathcal{P}$ is invertible.
@@
Let $\{s^{[q]}\}_{q=1}^{Q}$ be nonnegative normalized functions and $\{b_{d}\}_{d=1}^{D}$ be nonnegative bounded functions. Then for each $q$, the functional in 7 can be represented as a mapping $K^{[q]}:\mathcal{F}^{D}\rightarrow\mathcal{G}$ and is Gâteaux differentiable, and its Gâteaux derivative at $\boldsymbol{f}\in\mathcal{F}^{D}$ for $\boldsymbol{h}=[h_{1},\ldots,h_{D}]^{\boldsymbol{\mathsf{T}}}\in\mathcal{F}^{D}$ is computed as
$\mathcal{D}K^{[q]}(\boldsymbol{f},\boldsymbol{h})=-\int_{0}^{E_{\max}}\Phi^{[q]}(\boldsymbol{f};E)\left(\sum_{d=1}^{D}b_{d}(E)\mathcal{P}h_{d}\right)\mathrm{d}E,$ (10)
@@
$\Phi^{[q]}(\boldsymbol{f};E)=\frac{\displaystyle s^{[q]}(E)\exp\left(-\sum\limits_{d=1}^{D}b_{d}(E)\mathcal{P}f_{d}\right)}{\displaystyle\int_{0}^{E_{\max}}s^{[q]}(E)\exp\left(-\sum\limits_{d=1}^{D}b_{d}(E)\mathcal{P}f_{d}\right)\mathrm{d}E}.$ (11)
Moreover, both the $K^{[q]}$ and $\boldsymbol{K}:\mathcal{F}^{D}\rightarrow\mathcal{G}^{Q}$ are Fréchet differentiable.
@@
We start with analyzing the Gâteaux derivative of $K^{[q]}$ in 10. Notice that the $\Phi^{[q]}(\boldsymbol{f};E)$ in 11 is actually related to the geometric parameter $(\boldsymbol{\omega},\boldsymbol{x})\in\mathbb{T}^{n}$ since the term $\mathcal{P}f_{d}$ depends on that parameter as in 6. For a certain real constant set $\{C_{d}\}_{d=1}^{D}$, if there exists one $\hat{\boldsymbol{f}}=[\hat{f}_{1},\ldots,\hat{f}_{D}]^{\boldsymbol{\mathsf{T}}}$ such that
$\mathcal{P}\hat{f}_{d}(\boldsymbol{\omega},\boldsymbol{x})\equiv C_{d},\quad\forall(\boldsymbol{\omega},\boldsymbol{x})\in\mathbb{T}^{n},$ (12)
then $\Phi^{[q]}(\hat{\boldsymbol{f}};E)$ becomes to be independent of the geometric parameter $(\boldsymbol{\omega},\boldsymbol{x})$. Together with the linearity of $\mathcal{P}$, the Gâteaux derivative at $\hat{\boldsymbol{f}}$ for $\boldsymbol{h}$ can be rewritten as
$\mathcal{D}K^{[q]}(\hat{\boldsymbol{f}},\boldsymbol{h})=-\mathcal{P}\left(\sum_{d=1}^{D}\varphi^{[q]}_{d}h_{d}\right).$ (13)
@@
$\varphi^{[q]}_{d}=\int_{0}^{E_{\max}}\Phi^{[q]}(\hat{\boldsymbol{f}};E)b_{d}(E)\mathrm{d}E,$ (14)
@@
$\Phi^{[q]}(\hat{\boldsymbol{f}};E)=\frac{\displaystyle s^{[q]}(E)\exp\left(-\sum\limits_{d=1}^{D}b_{d}(E)C_{d}\right)}{\displaystyle\int_{0}^{E_{\max}}s^{[q]}(E)\exp\left(-\sum\limits_{d=1}^{D}b_{d}(E)C_{d}\right)\mathrm{d}E}.$
Obviously, perhaps the most simple example is taking $\hat{\boldsymbol{f}}=\boldsymbol{0}\in\mathcal{F}^{D}$ for $C_{d}=0$, $d=1,\ldots,D$ that satisfies the condition of 12. There may be other points that satisfy this condition, but searching for them is beyond the scope of the current research.
@@
$\mathcal{D}\boldsymbol{K}(\hat{\boldsymbol{f}},\cdot):\mathcal{F}^{D}$ $\longrightarrow\tilde{\mathcal{G}}^{Q}$
@@
[…]
@@
Assume that the conditions in proposition 1 hold, and the matrix $\boldsymbol{\varphi}$ is invertible. Then the $\mathcal{D}\boldsymbol{K}(\hat{\boldsymbol{f}},\cdot)$ in section 3.1.2 is invertible.

### v2

# AFIRE: Accurate and Fast Image Reconstruction Algorithm for Geometric-inconsistent Multispectral CT

@@ Abstract
For nonlinear multispectral computed tomography (CT), accurate and fast image reconstruction is challenging when the scanning geometries under different X-ray energy spectra are inconsistent or mismatched. Motivated by this, we propose an Accurate and Fast Image REconstruction (AFIRE) algorithm to address such problems in the case of mildly full scan. From the continuous (resp. discrete) setting, we discover that the derivative operator (gradient) of the involved nonlinear mapping at some special points, for example, at zero, can be represented as a composition (block multiplication) of a diagonal operator (matrix) composed of X-ray transforms (projection matrices) and a very small-scale matrix. Based on these insights, the AFIRE algorithm is proposed by leveraging the simplified Newton method. Under proper conditions, we establish the convergence theory of the proposed algorithm. Furthermore, numerical experiments are also carried out to verify that the proposed algorithm can accurately and effectively reconstruct the basis images in completely geometric-inconsistent dual-energy CT with noiseless and noisy projection data. Particularly, the proposed algorithm significantly outperforms some state-of-the-art methods in terms of accuracy and efficiency. Finally, the flexibility and extensibility of the proposed algorithm are also demonstrated.
@@
Multispectral CT, dual-energy CT, geometric inconsistency, accurate and fast image reconstruction algorithm, large-scale nonlinear inverse problem, one-step method, convergence analysis
@@ 1 Introduction
Multispectral CT (MSCT) is an imaging modality aiming to reconstruct material-specific (basis) images from multiple sets of attenuation data collected with different X-ray energy spectra. These basis images can subsequently be used to form virtual monochromatic images (VMIs) or the spatial distributions of linear attenuation coefficient (LAC) at energies of interest by linear combination [20, 4, 29]. MSCT has attracted increasing attention and research interest due to the development of hardware and the requirement of applications, as well as its advantages in considering the X-ray energy-dependent information over the conventional CT such as artifact suppression, contrast improvement and object separation [29]. Dual-energy CT (DECT), as the first generation of MSCT, has been implemented in clinical practice for about twenty years [19]. Despite the long history of DECT, its clinical and technical research is still in its infancy [21].
@@
It is well-known that the forward data model used in conventional CT assumes that all of the X-ray photons emitting from an X-ray source are monochromatic, namely, have the same energy [35]. This assumption makes the model linear, which causes the image reconstruction to become a linear inverse problem, but leading to difficulties in artifact suppression and limitations in material differentiation. In real-world CT, however, this assumption is not realistic because the output X-ray beams are polychromatic, consisting of photons with a wide range of energies. The distribution of photon energies is referred to as the energy spectrum. The LAC of the scanned material depends not only on its own elemental composition but also on the energy of the probing X-ray photons. Hence, the forward data model of MSCT may be described as a nonlinear mapping, and the corresponding image reconstruction can be formulated as a nonlinear inverse problem [4, 28, 31, 21, 29, 23].
@@
Although development of algorithms for reconstructing basis images in MSCT constitutes an important topic of active research, accurate and fast image reconstruction remains a challenging task, largely because its forward data model is nonlinear. In this work, we understand different reconstruction algorithms from the perspective of scanning geometric parameters under different X-ray energy spectra.
@@
One situation is that the scanning geometric parameters are consistent or matched under different energy spectra. That is to say, for each X-ray path, under every energy spectrum, there is always one X-ray beam passing through it. Here this case is called geometric-consistent MSCT, in which the image reconstruction can be accurately realized by a two-step method: step (i) estimates the basis sinograms from measurements through solving a lot of very small-scale independent systems of nonlinear equations along different X-ray paths, for instance, using Newton’s method, whereas step (ii) reconstructs basis images from the basis sinograms estimated, such as using the filtered back-projection (FBP) method. This method is commonly referred to as two-step data-domain-decomposition (DDD) method. As step (ii) involves the inversion of a linear system as in conventional CT, step (i) constitutes the key step of the DDD method in which a nonlinear data model needs to be inverted. Although many algorithms can invert the nonlinear model accurately, analysis of its solution existence, uniqueness and stability remains difficult. In recent years, several works studied the solution properties such as [25, 3, 5, 13, 18].
@@
The other is that the scanning geometric parameters are inconsistent or mismatched under different energy spectra. Namely, there is at least one X-ray path that all the X-ray beams under some energy spectrum do not pass through it. We call this case the geometric-inconsistent MSCT. A more extreme example is that for every X-ray path, there is always an energy spectrum under which no X-ray beams pass through it, which can be seen as completely geometric-inconsistent MSCT. The configurations of many commercial devices fall into this situation, such as the fast kV-switching DECT of GE [29].
@@
The development of accurate and fast reconstruction algorithms for this case is more challenging than the geometric-consistent case. The authors in [38] matched the originally mismatched neighboring projections of high- and low-energy spectra in DECT, and then proposed a two-step DDD method to invert those one-view-deviated projection pairs. Similarly, using the matched projections as in [38], an iterative FBP (IFBP) reconstruction method was proposed in [26], which applies a single Newton iteration and FBP alternately. However, due to the different geometric parameters of the matched projections, these approaches cannot result in an accurate image reconstruction.
In the last two decades, progress has been made in developing one-step methods for the image reconstruction in MSCT. To reconstruct the basis images directly, such kind of method often needs to devise an optimization model from a theoretical perspective, such as Bayesian inference and variational regularization, and then develop an iterative scheme to solve that model. For instance, the authors in [15, 8, 36, 27, 30] employed the maximum a posteriori estimation within Bayesian inference to formulate the optimization models, and the authors in [6, 9, 10, 17] constructed the optimization models from the perspective of variational regularization. Because of the nonlinear data model in MSCT, the optimization models designed are generally nonconvex. While these methods can often achieve high image quality through a large number of iterations, they are time-consuming, and their convergence analyses were mostly not considered. In addition, many studies also consider the problem as solving a large-scale system of nonlinear equations [37, 16].
@@
In this article, we aim to develop an image reconstruction algorithm of the one-step type that takes into account both accuracy and efficiency, has mathematical guarantee, and is suitable for the geometric-inconsistent MSCT.
@@
The rest of this paper is organized as follows. In section 2, we introduce the forward data models of MSCT. In section 3, we propose a novel algorithm to solve the image reconstruction problem in geometric-inconsistent MSCT. We then prove the convergence of the proposed algorithm in section 4. In section 5, numerical experiments are carried out to illustrate the performance of the proposed algorithm. The flexibility and extensibility of the proposed algorithm are discussed in section 6. Finally, we conclude the paper in section 7.
@@ 2.1 CC-data model
and $\mu(\boldsymbol{y},E)$ represents the energy-dependent LAC of interest at spatial position $\boldsymbol{y}\in\mathbb{R}^{n}$ ($n=2$ or 333) and energy $E\in[0,E_{\text{max}}]$, and $\mathbb{R}^{n}$ the real Euclidean space, $E_{\text{max}}$ the maximum energy in the scan [7, 21, 2]. The function $\mu$ of LAC is decomposed often into a linear combination of basis images $\{f_{d}\}_{d=1}^{D}$ as
@@
where $b_{d}(E)$ denotes the expansion coefficient, and $D$ the number of basis images [4, 21]. For example, for the typical X-ray energy range in diagnostic DECT, $D=2$ is chosen often as photoelectric effect and Compton scatter are the dominant factors contributing to $\mu$ in the physical effect decomposition model. For the basis material decomposition model, $b_{d}$ denotes the known function of mass attenuation coefficient (MAC) of the $d$-th basis material (e.g., water and bone); and $f_{d}$ represents the material-equivalent density function of the $d$-th basis material. Importantly, these two models are physically equivalent [21]. Consequently, using two basis materials is sufficient to accurately describe the linear attenuation characteristics.
@@
The formula in 3 is referred to as the CC-data model because $\theta^{[q]}$, $E$, and $\boldsymbol{y}$ are continuous variables [9].
@@ 2.2 DD-data model
In practice, data can be detected only for discrete energies and rays, and an image is reconstructed only on an array of discrete pixels/voxels. Hence, it is possible to develop a DD-data model, which may be constructed by implementing a specific discrete scheme into the CC-data model in 3 [9]. However, in general, there is no direct connection between the two models.
Assume that the energy domain $[0,E_{\text{max}}]$ is divided into $M$ intervals with equal size $\Delta_{E}$; the data $g_{j}^{[q]}$ are detected for discrete directions $\theta_{j}^{[q]}$ with $j=1,\dots,J_{q}$ under the $q$-th spectrum; each basis image $f_{d}$ is discretized by an array of size $I$. Then a DD-data model is obtained as
$g_{j}^{[q]}=\ln\sum_{m=1}^{M}s^{[q]}_{m}\exp\left(-\sum_{d=1}^{D}b_{dm}\sum_{i=1}^{I}p^{[q]}_{ji}f_{di}\right)\quad\text{for $j=1,\dots,J_{q}$,}$ (4)
@@ 3.1 The continuous setting
Before that, we introduce the required notation. Let $\mathcal{X}_{1},\ldots,\mathcal{X}_{D}$ be normed vector spaces with norms $\|\cdot\|_{\mathcal{X}_{1}},\ldots,\|\cdot\|_{\mathcal{X}_{D}}$, respectively. Their Cartesian product space is given by
[…]
The norm of an element $\boldsymbol{u}=[u_{1},\ldots,u_{D}]^{\boldsymbol{\mathsf{T}}}$ in this Cartesian product space can be defined as
$\|\boldsymbol{u}\|_{\mathbfcal{X}^{D}}:=\left(\sum_{d=1}^{D}\|u_{d}\|^{2}_{\mathcal{X}_{d}}\right)^{1/2}.$
Let $H^{\alpha}(\mathbb{R}^{n})$ be the Sobolev space of order $\alpha$ on $\mathbb{R}^{n}$, and $\Omega$ be a sufficiently regular open bounded subset in $\mathbb{R}^{n}$. Denote $H_{0}^{\alpha}(\Omega)=\{u\in H^{\alpha}(\mathbb{R}^{n})|\text{supp}(u)\subset\Omega\}$, where $\text{supp}(u):=\overline{\{\boldsymbol{y}\in\mathbb{R}^{n}|u(\boldsymbol{y})\neq 0\}}$ is the support of $u$.
@@ 3.1.1 Continuous MSCT image reconstruction problem
For simplicity, define $\mathbb{T}^{n}_{q}:=\{(\boldsymbol{\omega},\boldsymbol{x})|\boldsymbol{\omega}\in\mathbb{S}^{n-1}_{q}\subseteq\mathbb{S}^{n-1},\boldsymbol{x}\in\Omega_{q}\subseteq\boldsymbol{\omega}^{\bot}\}\subseteq\{(\boldsymbol{\omega},\boldsymbol{x})|\boldsymbol{\omega}\in\mathbb{S}^{n-1},\boldsymbol{x}\in\boldsymbol{\omega}^{\bot}\}=:\mathbb{T}^{n}$ as the set of scanning geometric parameters under the $q$-th energy spectrum, where $q=1,\ldots,Q$. Here $\mathbb{S}^{n-1}$ represents the unit sphere in $\mathbb{R}^{n}$ and $\boldsymbol{\omega}^{\bot}$ denotes the hyperplane perpendicular to $\boldsymbol{\omega}$.
Let
$\mathcal{P}_{q}f(\boldsymbol{\omega},\boldsymbol{x}):=\int_{-\infty}^{+\infty}f(\boldsymbol{x}+t\boldsymbol{\omega})~\mathrm{d}t,\quad(\boldsymbol{\omega},\boldsymbol{x})\in\mathbb{T}^{n}_{q},$ (5)
which is the X-ray transform of a certain function $f$ defined on $\mathbb{R}^{n}$, equivalently the Radon transform if $n=2$ [31]. Note that if $\mathbb{T}^{n}_{1}=\cdots=\mathbb{T}^{n}_{Q}$, then $\mathcal{P}_{1}=\cdots=\mathcal{P}_{Q}$, this situation is called geometric consistency, otherwise it is called geometric inconsistency.
For each energy spectrum $q$, denote the X-ray transform in 5 by the mapping
$\mathcal{P}_{q}:\mathcal{F}=H_{0}^{-1/2}(\Omega)\longrightarrow\tilde{\mathcal{G}}_{q}=\mathcal{P}_{q}(\mathcal{F})\subseteq\mathcal{L}^{2}(\mathbb{T}^{n}_{q})=\mathcal{G}_{q}.$
As the X-ray $L(\theta^{[q]})$ in 3 can be parametrized by $(\boldsymbol{\omega},\boldsymbol{x})$, we translate 5 into
$\mathcal{P}_{q}f(\boldsymbol{\omega},\boldsymbol{x})=\int_{L(\theta^{[q]})}f(\boldsymbol{y})~\mathrm{d}l.$ (6)
@@
$K^{[q]}(\boldsymbol{f}):=\ln\int_{0}^{E_{\max}}s^{[q]}(E)\exp\left(-\sum_{d=1}^{D}b_{d}(E)\mathcal{P}_{q}f_{d}\right)\mathrm{d}E,$ (7)
@@
Consequently, the aim of the continuous MSCT image reconstruction is to reconstruct $\boldsymbol{f}$ from the measured data $\boldsymbol{g}=[g^{[1]},\ldots,g^{[Q]}]^{\boldsymbol{\mathsf{T}}}$ such that
@@
where $\boldsymbol{K}(\boldsymbol{f})=[K^{[1]}(\boldsymbol{f}),\ldots,K^{[Q]}(\boldsymbol{f})]^{\boldsymbol{\mathsf{T}}}$.
@@ 3.1.2 The proposed algorithm for the continuous problem
[…]
$\mathcal{D}K^{[q]}(\boldsymbol{f},\boldsymbol{h})=-\int_{0}^{E_{\max}}\Phi^{[q]}(\boldsymbol{f};E)\left(\sum_{d=1}^{D}b_{d}(E)\mathcal{P}_{q}h_{d}\right)\mathrm{d}E,$ (10)
@@
$\Phi^{[q]}(\boldsymbol{f};E)=\frac{\displaystyle s^{[q]}(E)\exp\left(-\sum\limits_{d=1}^{D}b_{d}(E)\mathcal{P}_{q}f_{d}\right)}{\displaystyle\int_{0}^{E_{\max}}s^{[q]}(E)\exp\left(-\sum\limits_{d=1}^{D}b_{d}(E)\mathcal{P}_{q}f_{d}\right)\mathrm{d}E}.$ (11)
Moreover, both the $K^{[q]}$ and $\boldsymbol{K}:\mathbfcal{F}^{D}\rightarrow\mathbfcal{G}^{Q}$ are Fréchet differentiable.
@@
We start with analyzing the Gâteaux derivative of $K^{[q]}$ in 10. Notice that the $\Phi^{[q]}(\boldsymbol{f};E)$ in 11 is actually related to the geometric parameter $(\boldsymbol{\omega},\boldsymbol{x})\in\mathbb{T}^{n}_{q}$ since the term $\mathcal{P}_{q}f_{d}$ depends on that parameter as in 6. For a certain real constant set $C_{d}^{[q]}\}_{d=1}^{D}$, if there exists one $\hat{\boldsymbol{f}}=[\hat{f}_{1},\ldots,\hat{f}_{D}]^{\boldsymbol{\mathsf{T}}}\in\mathbfcal{F}^{D}$ such that
$\mathcal{P}_{q}\hat{f}_{d}(\boldsymbol{\omega},\boldsymbol{x})\equiv C_{d}^{[q]},\quad\forall(\boldsymbol{\omega},\boldsymbol{x})\in\mathbb{T}^{n}_{q},$ (12)
then $\Phi^{[q]}(\hat{\boldsymbol{f}};E)$ is independent of the geometric parameter $(\boldsymbol{\omega},\boldsymbol{x})$. Together with the linearity of $\mathcal{P}_{q}$, the Gâteaux derivative at $\hat{\boldsymbol{f}}$ for $\boldsymbol{h}$ can be rewritten as
$\mathcal{D}K^{[q]}(\hat{\boldsymbol{f}},\boldsymbol{h})=-\mathcal{P}_{q}\left(\sum_{d=1}^{D}\varphi^{[q]}_{d}h_{d}\right).$ (13)
@@
$\varphi^{[q]}_{d}=\int_{0}^{E_{\max}}\Phi^{[q]}(\hat{\boldsymbol{f}};E)b_{d}(E)~\mathrm{d}E,$ (14)
@@
$\Phi^{[q]}(\hat{\boldsymbol{f}};E)=\frac{\displaystyle s^{[q]}(E)\exp\left(-\sum\limits_{d=1}^{D}b_{d}(E)C_{d}^{[q]}\right)}{\displaystyle\int_{0}^{E_{\max}}s^{[q]}(E)\exp\left(-\sum\limits_{d=1}^{D}b_{d}(E)C_{d}^{[q]}\right)\mathrm{d}E}.$
Remark 1.
The simplest example is taking $\hat{\boldsymbol{f}}=\boldsymbol{0}\in\mathbfcal{F}^{D}$ for all of the $C_{d}^{[q]}=0$, $d=1,\ldots,D$ and $q=1,\ldots,Q$ that satisfies the condition of 12. Although there may be other points that satisfy this condition, their explicit identification is not critical and is beyond the scope of the current research. This is because both the practical implementation and the theoretical convergence of the proposed algorithm are independent of the explicit form of such a point.
@@
$\mathcal{D}\boldsymbol{K}(\hat{\boldsymbol{f}},\cdot):\mathbfcal{F}^{D}$ $\longrightarrow\tilde{\mathbfcal{G}}^{Q}$
@@
where $\boldsymbol{\mathcal{P}}=\operatorname*{\texttt{diag}}(\mathcal{P}_{1},\ldots,\mathcal{P}_{Q})$ denotes a diagonal operator composed of $Q$ X-ray transforms, and $\boldsymbol{\varphi}=\bigl{(}\varphi^{[q]}_{d}\bigr{)}_{Q\times D}$ represents the constant matrix with entries $\varphi^{[q]}_{d}$ given in 14.
@@
Assume that $Q=D$, the conditions in proposition 1 hold, and the X-ray transform $\mathcal{P}_{q}:\mathcal{F}\rightarrow\tilde{\mathcal{G}}_{q}$ for each $q$ and the matrix $\boldsymbol{\varphi}$ are invertible. Then the $\mathcal{D}\boldsymbol{K}(\hat{\boldsymbol{f}},\cdot)$ in section 3.1.2 is invertible.